OpenStudy (anonymous):

write an equation of the line that passes through (-8,8) and (1,3)

OpenStudy (anonymous):

Equation of a line, y = mx+b

OpenStudy (anonymous):

Find the slope first, m = slope, m = y2-y1/(x2-x1), I gtg now hope that helps, bye!

OpenStudy (anonymous):

i get that far then i end up with wrong answer still

OpenStudy (anonymous):

y=mx+b find the slope m=y2-y1/x2-x1 m=3-8/1+8 m=-5/9 so far we have y=-5/9x+b now plug in any point. i'll use (1,3). x=1, y=3 3=-5(1)/9+b 3=-5/9+b 27=-5+9b 32=9b b=32/9 so y=mx+b y=(-5/9)x+32/9